### v2

@@ Acknowledgments
For the muon simulations [29] the expected flux and muon spectrum at an overburden of 7.4 m w.e. were calculated from those at Earth’s surface ([42],[43]) and propagated through the shield. The resulting spectrum was validated by comparing it to the Conus+ data without the applied muon veto cut and good agreement was found. The muon veto cut was then applied by multiplying the simulation output by a factor of 0.01 for all energies greater than 2 keVee, corresponding to a muon veto efficiency of 99% for these energy regions.
For energies below 2 keVee a different approach was taken. Here, simulations show an inefficiency in the tagging ability of the muon veto system due to the setup of the shield. As shown in figure 1, the outer muon veto is located under a layer of lead in the Conus+ shield. As such, it is possible that muons pass through this outermost lead layer without hitting one of the muon veto layers. These muons can induce electromagnetic showers in the outer lead layer, which propagate through the shield and are registered in the Germanium detectors. However, because no muon passes through any muon veto layer in such an event, the energy deposition in the plastic scintillator plates will be much lower, resulting in a greatly reduced tagging efficiency of these events. Simulations show that at very low energies (below 0.4 keVee) up to 80% of all muon-induced background come from such events. Based on this simulation output, this inefficiency was modeled with a polynomial and accounted for in the final muon veto efficiency below 2 keVee. The resulting efficiency drops towards lower energies with its minimum being 97% for energies below 0.4 keVee.
Using this approach, the overall background contribution of cosmic ray muons was found to be counts d-1kg-1 in each detector in the energy range between 0.4 keVee and 1 keVee, which corresponds to approximately a third of the background counts in this region, respectively.
@@
The remaining background in each detector is made up of many different components similar to the situation in Brokdorf (KBR) [29]. There are no hints that the detector upgrades or the movement of the setup were introducing any contamination. The background model includes cosmogenically induced isotopes in the Copper parts (57Co, 60Co, 54Mn) of the cryostat and the Germanium crystals (57Co, 68Ge, 68Ga, 65Zn, 3H), Radon inside the detector chamber, 210Pb inside the cryostat and shield, metastable germanium states (71mGe, 73mGe, 75mGe) and inert gases coming from the reactor (85Kr, 135Xe, 3H). The results of the simulation of these components were scaled to be in accordance with either the rates of gamma lines produced by the decay of these isotopes in the spectrum (e.g. for radon) or with screening measurements done prior to the installation. The listed contributions are typically very subdominant in the region of interest, with the decay of radon inside the detector chamber being the only exception. This contribution results in a background rate of counts d-1kg-1 between 0.4 keVee and 1 keVee for the C5 detector (C2: counts d-1kg-1; C3: counts d-1kg-1). The difference in radon contributions is likely due to the distribution of the radon decay products inside the detector chamber. As such, the C5 detector has the lowest radon component due to its proximity to the entry point of the radon free air. The listed values correspond to approximately 5% of the background in this energy region. Furthermore, radon decay has a high impact on energies above 100 keVee, where it can contribute up to 60% of the measured background. Slow pulses, arising from decays on the surface of the diode are included in the model [29].
Leakage test component
As previously described in [27], during the final run of the Conus experiment in Brokdorf an additional background component had to be introduced in the background model. This component was present after ventilation of the cryostats with argon gas to avoid mechanical deformation during a regular leakage test at KBR in July 2019 [29]. Of the four detectors that were used in the Conus experiment and were therefore affected by this additional background, two (C2 and C3) are used for the analysis presented in this work. Indeed, the simulations and the background model show that an additional component with the same shape is still present in the background of these two detectors but is absent in C5 which was not at KBR at the time. This additional background is constant during reactor on and off periods. Therefore, the leakage test component was again included in the background model of C2 and C3 by modeling it using a function with two parameters, as in [27]. The resulting impact is below 10% in the energy region between 0.4 keVee and 1 keVee.
@@
In addition to the background sources listed above, the background model of Conus+ also accounts for experimental differences during data collection with the reactor off. The first of these differences is induced by the fact that during a reactor outage the head of the drywell, the containment structure immediately surrounding the core of the reactor, is placed directly above the Conus+ room. This drywell head is made of 3.8 cm of steel and therefore increases the overburden of the experiment by approximately 0.3 m w.e. which results in a reduction of 19% in the flux of cosmogenic neutrons and a reduction of 3% in the cosmic ray muon flux. As a result, these two background contributions are reduced accordingly.
A second difference in the background model of the reactor off period comes from a more effective radon removal in the detector chamber. Over the course of run 1 shield tightness and flushing with radon free air were improved. As a result, the radon contribution in reactor off time is reduced by approximately a factor of 4 to 6 compared to reactor on. The radon contribution in the reactor off data was again scaled to match the count rate in the gamma lines induced by the radon decay.
In addition, reactor-correlated background components, such as reactor neutrons and high energy gammas, e.g. from 16N, were investigated. Their impact was found to be negligible in all energy regions, including the region of interest. The resulting background model for all three detectors can be seen in figures 2 and figure 6.
In figure 7, the reactor on and off data sets are compared with the background models. If the calculated difference between on and off phases is added to the reactor off data, good agreement with the measured on data is found above the signal region. For all three detectors, good agreement was found between models and data for both reactor on (above 0.4 keVee) and reactor off phases (full spectrum). Since the data in the reactor off phase is statistically limited, the background model still has an important role in the analysis.
Table 3 gives an overview of the main background sources and their contribution in the energy range between 0.4 keVee and 1 keVee. A dedicated publication focusing on the full background decomposition is planned for the near future.
[0.4-1.0] keVee | ||||||
Detector | C5 | C2 | C3 | |||
Rector period | On | Off | On | Off | On | Off |
Cosmogenic neutrons | 21.6 3.1 | 17.7 2.5 | 21.6 3.1 | 17.7 2.5 | 21.6 3.1 | 17.7 2.5 |
Cosmogenic muons | 17.4 0.3 | 16.9 0.3 | 17.4 0.3 | 16.9 0.3 | 17.4 0.3 | 16.9 0.3 |
Radon | 1.9 0.1 | 0.3 0.1 | 2.8 0.1 | 0.7 0.1 | 2.6 0.1 | 0.7 0.1 |
Other | 2.0 0.2 | 1.2 0.2 | 6.4 0.5 | 5.6 0.5 | 5.6 0.5 | 4.8 0.5 |
Leakage test component | - | - | 3.0 0.5 | 3.0 0.5 | 0.8 0.2 | 0.8 0.2 |
Total (Model) | 42.9 3.1 | 35.8 2.5 | 52.3 3.3 | 45.1 2.7 | 49.3 3.1 | 42.2 2.7 |
Total (Data) | 43.5 1.1 | 33.4 1.8 | 50.7 1.2 | 45.3 1.3 | 48.8 1.2 | 42.5 2.0 |
@@
The ratio of the ionization energy released by nuclear recoil in a CENS event and the ionization energy of electrons of the same energy is given by the quenching factor. In the Conus+ analysis, the energy dependent signal quenching is described by the Lindhard model [32] with a quenching parameter as determined in [34]. Alternative quenching descriptions are also tested as described in the supplemental material of [27] including a linear and cubic functional form to describe the increased quenching factor compared to Lindhard theory found in [33].
The signal predictions for the description of these different quenching models are shown in figure 8 together with the difference between the data in the reactor on phase and the background model. A signal rate of () events is expected for the linear function, while for the cubic function () events are predicted. Both numbers are significantly higher than the neutrino rate extracted from the Conus+ data. The standard Lindhard model clearly provides the best description of the reactor correlated excess at low energy. In figure 8 we have also included an additional data point at 200 eVee. This one is obtained using the information shown in figure 4 of [17]. There, a CENS signal was approximated with an exponential using two parameters, the amplitude at 200 eVee (A0.2) and a decay constant. The favored value for A0.2 is shown within a 2 contour in a 2-dimensional plot of the two parameters. We scaled this A0.2 value to our exposure and corrected for the difference in neutrino flux. The value obtained was added in figure 8. As expected, it matches the description of Lindhard with a linear increase added at low energies but is in clear conflict with the Conus+ data. The lower error bar corresponds to the smallest value of A0.2 in the 2 contour. This value is also clearly ruled out by the Conus+ data points.
@@
and are the likelihood functions for reactor on and off periods. Gaussian pull terms for the systematic uncertainties are also included. Here, the first term represents the pull terms for correlated parameters, namely the trigger efficiency parameters, and the second term represents pull terms for uncorrelated parameters, like the active mass of the detectors, the reactor neutrino flux and the uncertainty on the energy scale calibration of the spectra. Parameters that were experimentally determined are pulled to their measured values. A background scaling factor is also included.
Detector | [eVee] | mass [kg] | live time | Signal events data | Signal predicted | Ratio |
C2 | 180 | 117 days | ||||
C3 | 160 | 110 days | ||||
C5 | 170 | 119 days | ||||
combined | ||||||
[…]
@@
Fits using just single detectors independently give consistent results as shown in Table 4. To cross-check the result, fits were performed by two independent likelihoods with different approaches concerning the uncertainty on the quenching model. Likelihood A applied a predicted CENS spectrum with a fixed k parameter () while introducing a fourth order polynomial to vary the shape of the signal spectrum applying Gaussian pull terms on each parameter. In likelihood B, the k value of the Lindhard model is a fit parameter with a pull term. Moreover, there are some differences in the treatment of the non-linearity corrections and the minimization algorithms between the two likelihood fits. In the end, the results of both likelihood implementations agree within 2%.
The overview of the systematic uncertainties in signal prediction is summarized in table 5. To be conservative, we assume that all of them are fully correlated between single detectors. Currently, it is dominated by the 5 eVee uncertainty in the energy calibration. Future extended calibrations will allow us to improve this value to 3 eVee, reducing the contribution of this systematic term. To achieve a total signal prediction uncertainty below 8% new improved quenching factor measurements would be needed.
Prediction uncertainties | |
Uncertainty | Contribution |
Energy threshold | 14.1% |
Quenching Ge | 7.3% |
Reactor neutrino flux | 4.6% |
Cross-section | 3.2% |
Active mass Ge | 1.1% |
Trigger efficiency | 0.7% |
All combined | 17% |
CENS result uncertainties | |
Uncertainty | Contribution |
Likelihood fit | 86 |
Fit method | 7 |
Background model | 40 |
Non-linearity implementation | 47 |
All combined | 106 |
The likelihood fit itself gives a result of () CENS counts in the combined fit of the three detectors. Additional systematic uncertainties not implemented as pull terms are estimated in a second step and added in quadrature, giving the final uncertainty of signal counts. The different contributions are listed in table 5. The biggest contribution comes from the likelihood fit, combining both statistic and systematic uncertainties considering the pull terms in the likelihood function. The second biggest impact is the uncertainty due to the implementation of the non-linearity correction, explained above. This is caused by the choice of a specific function to describe and fit the deviations from a purely linear energy scale in figure 5. This uncertainty was obtained through a systematic study of other possible fit functions and checking the impact on the likelihood result. An additional uncertainty from the background model description originates mainly from uncertainties on the muon and neutron flux in the Conus+ room as well as the leakage component. The uncertainty was obtained by varying the background model within the model constraints and subsequently performing the likelihood fits. Lastly, the fit method itself, i.e. the difference between Likelihood A and B, as mentioned above, has a small additional impact on the overall uncertainty.
@@
The fractional contribution from antineutrinos of different energies to the signal expectation is quantified in figure 9. At the current detector threshold, we are sensitive to antineutrino energies above 5 MeV. With lower detection threshold, the steep rise in the antineutrino flux towards low energies will result in significantly higher signal expectations [28].

### v3

@@ Acknowledgments
For the muon simulations [29] the expected flux and muon spectrum at an overburden of 7.4 m w.e. were calculated from those at Earth’s surface ([42],[43]) and propagated through the shield. The resulting spectrum was validated by comparing it with the Conus+ data without the applied muon veto cut. Good agreement was found. The muon veto cut was then applied by multiplying the simulation output by a factor of 0.01 for all energies greater than 2 keVee, corresponding to a muon veto efficiency of 99%.
For energies below 2 keVee a different approach was taken. Here, simulations show an inefficiency in the tagging ability of the muon veto system due to the setup of the shield. The outer muon veto is located under a layer of lead in the Conus+ shield. As such, it is possible for muons to pass through this outermost lead layer without hitting one of the muon veto layers. These muons can induce electromagnetic showers in the outer lead layer, which propagate through the shield and are registered in the Germanium detectors. However, because no muon passes through any muon veto layer in such an event, the energy deposition in the plastic scintillator plates will be much lower, resulting in a greatly reduced tagging efficiency of these events. Simulations show that at energies below 0.4 keVee up to 80% of all muon-induced background come from such events. Based on this simulation output, this inefficiency was modeled with a polynomial and accounted for in the final muon veto efficiency. The resulting efficiency drops towards lower energies with its minimum being 97% below 0.4 keVee. Using this approach, the overall background contribution of cosmic ray muons was found to be counts d-1kg-1 in each detector in the energy range between 0.4 keVee and 1 keVee, which corresponds to approximately a third of the background counts in this region.
Leakage test component
During the final run of the Conus experiment in at KBR an additional background component had to be introduced in the background model [27]. This component was present after ventilation of the cryostats with argon gas to avoid mechanical deformation during a regular leakage test at KBR in July 2019 [29]. Of the four detectors used in the Conus experiment, two (C2 and C3) are used for the analysis presented in this work and are affected by this additional background. Indeed, the simulations and the background model show that an additional component with the same shape is still present in the background of these two detectors but is absent in C5 which was not at KBR at the time. This additional background is constant during reactor on and off periods. Therefore, the leakage test component was again included in the background model of C2 and C3 by modeling it using a function with two parameters, as in [27]. The resulting impact is below 10% in the energy region between 0.4 keVee and 1 keVee.
@@
The remaining background in each detector is made up of many different components, similar to the situation in KBR [29]. There are no hints that the detector upgrades or the movement of the setup were introducing any contamination. The background model includes cosmogenically induced isotopes in the Copper parts (57Co, 60Co, 54Mn) of the cryostat and the Germanium crystals (57Co, 68Ge, 68Ga, 65Zn, 3H), Radon inside the detector chamber, 210Pb inside the cryostat and shield, metastable germanium states (71mGe, 73mGe, 75mGe) and inert gases coming from the reactor (85Kr, 135Xe, 3H). The results of the simulation of these components were scaled to be in accordance with either the rates of gamma lines produced by the decay of these isotopes in the spectrum (e.g. for radon) or with screening measurements performed prior to the installation. The listed contributions are typically very subdominant in the region of interest, with the decay of radon inside the detector chamber being the only exception. This contribution results in a background rate of counts d-1kg-1 between 0.4 keVee and 1 keVee for the C5 detector (C2: counts d-1kg-1; C3: counts d-1kg-1). These values correspond to approximately 5% of the background in this energy region. Radon decays have a high impact on energies above 100 keVee, where they can contribute up to 60% of the measured background. Slow pulses arising from decays on the surface of the diode and in the transition layer are included in the model [29]. Energy depositions from these events can be stopped within the transition layer and the released charge diffuses slowly into the active volume, resulting in long rise times and incomplete charge collection. Their impact is accounted for by registering the exact coordinates of an interaction in the germanium crystal. If the coordinates place it within the transition layer of the crystal, the energy of the event is shifted towards lower energies using a sigmoid-like function. Details on this procedure can be found in [26].
@@
The background model accounts for the experimental differences during data collection with the reactor off. The first of these differences is induced by the fact that during a reactor outage, the drywell head from the containment structure surrounding the reactor core is placed directly above the Conus+ room. This drywell head is made of 3.8 cm steel and therefore increases the overburden of the experiment by approximately 0.3 m w.e. which results in a reduction of 19% in the flux of cosmogenic neutrons and a reduction of 3% in the flux of cosmic ray muons. As a result, these two background contributions are reduced accordingly. The second difference in the background model of the reactor off period comes from a more effective removal of radon in the detector chamber. During the course of run 1 shield tightness and radon free air flushing were improved. As a result, the radon contribution in reactor off time is reduced by approximately a factor of 4 to 6 compared to reactor on. The radon contributions are scaled to match the count rates in the gamma lines induced by the radon decay. In addition, reactor-correlated background components, such as reactor neutrons and high energy gammas, e.g. from 16N, were investigated. Their impact was found to be negligible in all energy regions, including the region of interest. The resulting background model for all three detectors can be seen in Fig. 2 and Extended Data Fig. 6. The time-dependent contributions of the M-shell line of the 68Ge/71Ge decays are included in the background model. The difference between the M-shell contribution in reactor on and off periods was estimated based on the measured K-shell count rates. This difference was found to be below 1% of the observed neutrino signal.
a
b
c
d
In Extended Data Fig. 7, the reactor on and off data sets are compared with the background models. If the calculated difference between on and off phases is added to the reactor off data, good agreement with the measured on data is found above the signal region. For all detectors background models and data are fully consistent above 0.4 keVee). Since the data in the reactor off phase is statistically limited, the background model still plays an important role in the analysis. Extended Data Fig. 8 shows the fractional contribution of the main components to the total background rate as predicted in the model.
a
b
@@
The ratio of the ionization energy released by nuclear recoil in a CENS event and the ionization energy of electrons of the same energy is given by the quenching factor. In the Conus+ analysis, the energy dependent signal quenching is described by the Lindhard model [32] with a quenching parameter as determined in [34]. Alternative quenching descriptions are also tested as described in the supplemental material of [27] including a linear and cubic functional form to describe the increased quenching factor compared to Lindhard theory found in [33]. It was shown that the Migdal effect is subdominant with respect to the CENS signal in our region of interest [36].
The signal predictions for the different quenching descriptions are shown in Extended Data Fig. 9 together with the difference between the data in the reactor on phase and the background model. A signal rate of () events is expected for the linear function, while for the cubic function () events are predicted. Both numbers are significantly higher than the neutrino rate extracted from the Conus+ data. The standard Lindhard model clearly provides the best description of the reactor-correlated excess at low energy. In Extended Data Fig. 9 we have included an additional data point at 200 eVee, which is obtained using the information shown in figure 4 of [17]. There, a CENS signal was approximated with an exponential using two parameters, the amplitude at 200 eVee (A0.2) and a decay constant. The favored value for A0.2 is shown within a 2 contour in a 2-dimensional plot of the two parameters. We scaled this A0.2 value to our exposure and corrected for the difference in neutrino flux. As expected, it matches the description of Lindhard with a linear increase added at low energies but is in clear conflict with the Conus+ data. The lower error bar corresponds to the smallest value of A0.2 in the 2 contour. This value is also clearly ruled out by the Conus+ data points.
@@
and are the binned likelihood functions for reactor on and off periods. Gaussian pull terms for the systematic uncertainties are also included. Here, the first term represents the pull terms for correlated parameters, namely the trigger efficiency parameters, and the second term represents pull terms for uncorrelated parameters, like the active mass of the detectors, the reactor neutrino flux and the uncertainty on the energy scale calibration of the spectra. Parameters that were experimentally determined are pulled to their measured values. The background scaling factor , an additional fit parameter for the overall normalization of the background model, is also included and pulled to one. The binned likelihood functions assume a Poisson distribution and have the form
where N is the number of bins in the region of interest, is the bin content in the histograms of the measured data and is the bin content of the model. For a single detector, the model for reactor on data is calculated from the sum of the background model (scaled with ) and the predicted CENS spectrum (scaled with the signal parameter ) by taking into account the live time of the experiment ( and ), the active volume of the detector (), the dead time correction (), the fission flux () and a multiplication factor (), summarizing uncertainties of the detector response. In total:
The combined fit minimizes and for all three detectors simultaneously while the signal parameter , which indicates the number of CENS counts, is shared among detectors.
@@
Fits using just single detectors independently give consistent results as shown in Extended Data Table 3. To cross-check the result, fits were performed by two independent likelihoods with different approaches concerning the uncertainty on the quenching model. Likelihood A applied a predicted CENS spectrum with a fixed k parameter () while introducing a fourth order polynomial to vary the shape of the signal spectrum applying Gaussian pull terms on each parameter. In likelihood B, the k value of the Lindhard model is a fit parameter with a pull term. Moreover, there are some differences in the treatment of the non-linearity corrections and the minimization algorithms between the two likelihood fits. The results of both likelihood implementations agree within 2%.
For the signal prediction, we choose the Helm parameterization of the nuclear form factor [44, 45] and a reactor antineutrino spectrum based on the data-driven approach of [31]. The given antineutrino spectra are adjusted to the fission fraction of the Leibstadt reactor and augmented with measured antineutrino spectra at energies above 8 MeV [46] and simulation data below 1.8 MeV (the threshold of inverse beta decay) [47]. We assume no uncertainty on the nuclear form factor, but account for a 2% uncertainty on the weak mixing angle [48] leading to an overall uncertainty of 3.2% on the CENS cross section. The shape of the applied antineutrino spectra, the thermal power of the reactor, fission fractions and energy releases per fission of the relevant fission isotopes contribute a combined 4.6% error to the expected event rate. Further uncertainties on quenching, detector active mass, trigger efficiency and the energy threshold are taken into account and listed in Extended Data Table 4. All contributions lead to an overall uncertainty of 17% on the signal prediction. To be conservative, we assume prediction uncertainties are fully correlated between single detectors. Currently, the dominant uncertainty is from the energy calibration. Future extended calibrations will allow us to reduce this uncertainty from 5 eVee to 3 eVee.
The likelihood fit itself gives a result of () CENS counts in the combined fit of the three detectors. Additional systematic uncertainties not implemented as pull terms are evaluated in a second step and added in quadrature, giving the final uncertainty of signal counts, as shown in the Extended Data Table 4. The non-linearity term was obtained by varying the correction parameters and checking the impact on the likelihood result. The uncertainties of the calibration points in Extended Data Fig. 5 were used to generate the parameter variations. A Gaussian fit was performed on the distribution of the central values of the new likelihood results, where its 1 value was taken as additional systematic uncertainty. The term related to the background model was studied in the same way by varying muon flux, neutron flux, and the leakage component. The variations were 14% (taken from [40]), 6% (based on [43]) and 10%, respectively. Additionally, the systematic uncertainty of the fit method was estimated by the difference between likelihood A and B.
@@
The fractional contribution from antineutrinos of different energies to the signal expectation is quantified in Extended Data Fig. 10. At the current detector threshold, we are sensitive to antineutrino energies above 5 MeV. With lower detection thresholds, the steep rise in antineutrino flux toward low energies will result in significantly higher signal expectations [28].